10 L of o2to gas to molecules


1
Expert's answer
2021-11-12T04:05:04-0500

At STP, 1 mole of gas occupies 22.4 l.

n (O2) = 10 / 22.4 = 0.45 mol

A mole (mol) is the amount of a substance that contains 6.02 × 10 23 representative particles of that substance.

N (O2) = 0.45 x 6.02 × 10 23 = 2.69 x 10 23 molecules.